OpenStudy (anonymous):

Find the equation of the line that contains the point whose coordinates are (6, -1) and has slope 2/3.

OpenStudy (anonymous):

use point-slope formula \[\large y - y1 = m(x - x1)\] where m = 2/3, y1=-1 and x1=6

OpenStudy (anonymous):

what is the equation?

OpenStudy (anonymous):

You tell me.

OpenStudy (anonymous):

I have no clue, i was absent that day and the book makes no sense

OpenStudy (anonymous):

plug in all the numbers i have labeled and solve for y.. ill give you a start.\[y - (-1) = \frac{2}{3}(x - 6)\] now all you do is solve for y.

OpenStudy (anonymous):

subsistute the values that tyler 1992 gave you into the formula, you will get your answer

OpenStudy (anonymous):

im so lost